# ChatReID: Open-ended Interactive Person Retrieval via Hierarchical Progressive Tuning for Vision Language Models

### v1

@@ 3.2.2 Data Engine
The current LLM model exhibits superior text-to-image retrieval capabilities compared to image-to-image retrieval in the person Re-ID task. Consequently, we augment the data volume for the image-to-image tasks to significantly enhance its matching ability during this training stage. The two image-to-image tasks comprise approximately 25% to 30% of the overall data volume, whereas the remaining tasks each constitute roughly 10%.
@@
Image-image data and text-to-image data construction. The quantity of sampled gallery images for the Many-to-Many Retrieval task adheres to a uniform distribution $N\sim\text{Uniform}(a,b)$. This primarily aims to complicate the learning process and to inhibit the model from excessively depending on selecting responses from a particular gallery size. The distribution of the response $P(n)$ is obtained based on inverse transform sampling. This approach allows us to balance the probability that each number falls within different intervals, ensuring that the probability of recording any number eventually is uniform across all numbers. The probability $P(n=k)$ is given by:
@@
This formulation ensures that each number in the range has an equal chance of being recorded, leading to a uniform coverage of all possible values.This primarily aims to prevent the model from acquiring patterns in the distribution of responses, hence diminishing the complexity of learning.
Attribute-based data construction. The SYNTH-PEDES dataset includes precise annotations of pedestrian attributes to assist with the attribute-based task. Initially, we clean the templated pedestrian attribute annotations to isolate just the properties that are present in the reported pedestrian. In the task of attribute-based generation of image comparison descriptions, we do a preliminary comparison of the sampled pedestrian attributes and utilize the results as ground truth input for the model.
3.2.3 Joint Training
We carry over the encoder from Stage 1 but replace the decoder to continue full-parameter training. The previous decoder’s strong captioning focus could hinder outcomes in this stage, so we use a freshly pre-trained decoder instead. During this stage, we adjust the size of input images to accommodate a large number of gallery images. Images larger than 384$\times$192 were rescaled to 384$\times$192, while images below this threshold were left unaltered.
3.3 Decoder Post-training
After compressing the general pedestrian discrimination information via the above two stages, ChatReID has foundational skills in person identity matching. During this stage, we fine-tune the language decoder to improve the practicality of GRP, deepening the model’s understanding of Re-ID objectives and further enhancing its instruction-following abilities for effective real-world deployment.
3.3.1 Data Engine
We rethink the assessment of the person Re-ID problem and formulate a question construction in the style of Visual Question Answering (VQA) utilizing a multi-modal large language model from a practical application perspective. The instruction tuning data amounts to 500K.
Person Re-ID. The most common type of query involves providing a text or image query alongside a gallery set, with the goal of identifying images in the gallery that match the query’s identity. However, for current LVLMs, retrieving an indefinite-length sequence of images from an indefinite gallery set based on specific criteria is highly challenging. To address this, we transform this complex retrieval task into a multi-turn best-choice problem through engineering enhancements. Specifically, at each step, the model selects the most similar image, making a locally optimal choice, which simplifies the problem and reduces computational complexity.
@@ 3.3.1 Data Engine
Adaptive Multi-Modal Person Retrieval. Finally, in practical scenarios, multi-turn retrieval based on limited query information often results in a set of responses with relatively low accuracy. As retrieval continues, users typically gather more query details to improve subsequent searches. For example, beginning with attribute-based retrieval, if the correct match appears among the initial results, the user can use that match as a new query to refine further results. Alternatively, users can start with image-based retrieval and add textual details to address complexities such as clothing changes or long-term recognition tasks. By integrating multi-modal information, this approach significantly improves both the flexibility and accuracy of pedestrian re-identification, advancing its real-world applicability. To support this, we design a training objective for handling mixed queries with multi-modal inputs.
3.3.2 Post-training
We inherit the full structure from Stage 2, training with the encoder frozen. With pedestrian recognition skills already developed, this stage focuses on enhancing the model’s instruction-following capabilities. The input image resolution is preserved without any adjustments.
@@ 4 Experiments
4.1 Experimental Setups
Datasets. We adhere to the datasets used in Stage 3, using only the test sets. We utilize the commonly used metrics in person Re-ID, such as Cumulative Match Characteristic (CMC) and Mean Average Precision (mAP). Results are presented in terms of mAP/Rank-1.
Evaluation strategy. Evaluating LVLMs for person Re-ID in a VQA format presents two main challenges. First, the large gallery sets in person Re-ID datasets exceed the token limits of current LVLMs. Second, retrieving an indefinite sequence of images based on specific criteria from a vast gallery is difficult for these models. To address these issues, we implement several engineering optimizations. We begin by training a base model to filter the gallery, calculating feature similarity between the query image and gallery images. Only images with similarity above a threshold $\tau$ are retained, significantly reducing gallery size and improving evaluation efficiency without affecting accuracy. Additionally, as described in Stage 3, we convert this retrieval task into a multi-turn best-choice problem. Finally, we sample non-overlapping images from the filtered gallery and pair each selected image with the query and rank them based on response confidence, producing the final similarity list.
@@ 4.1 Experimental Setups
Table 3: Comparison of ChatReID with SOTA methods across two CC Re-ID datasets.
@@ 4.2 Experimental Results
Standard Person Re-ID. As shown in Tab. 2, ChatReID achieves impressive results across all standard person Re-ID datasets, particularly excelling on challenging datasets. The most notable improvement is observed on the complex and large-scale MSMT17 dataset, where ChatReID achieves a 15.1% increase in mAP. This substantial gain highlights the strength of our three-stage tuning strategy, which effectively extracts robust pedestrian-specific features, thereby improving model resilience in demanding scenarios. On the simpler Market1501 dataset, the performance gain is smaller, with a 2.9% increase in mAP. This modest improvement is expected, as existing methods have already achieved high accuracy on this dataset. Nonetheless, ChatReID’s measurable boost in performance demonstrates the consistent effectiveness of our approach, even when applied to datasets with strong baseline results.
CC Re-ID. The results in Tab. 3 indicate that ChatReID demonstrates significant mAP enhancements over IRM on both the LTCC and PRCC datasets, with increases of 24.3% and 23.3% in mAP, respectively. Although both IRM and ChatReID incorporate several person Re-ID settings, ChatReID outperforms for two primary reasons. First, all three stages of our tuning are based on image-text pairs, where textual information aids the visual encoder in learning variations in clothing more effectively. Second, we introduced innovative attribute-based training tasks in Stage 2, which enhance the model’s ability to recognize cloth-changing scenarios.
Table 4: Comparison of ChatReID with SOTA methods across two VI-ReID datasets.
@@
Table 5: Ablation Study. Performance comparison of different stage combinations in our hierarchical progressive learning strategy across ten benchmarks.
@@
Market1501 MSMT17 CUHK03 PRCC LTCC SYSU-MM01 RegDB CUHK-PEDES. ICFG-PEDES RSTPReid
@@
VI-ReID. As shown in Tab. 4, ChatReID also demonstrates significant improvements on VI-ReID datasets. Specifically, it achieves an 11.8% mAP increase on SYSU-MM01 and a 10.7% mAP increase on RegDB. These gains are primarily due to our approach in Stage 2, where we treated the VI-ReID task as an image-image matching task and slightly increased the proportion of VI data in the tuning process. This adjustment allowed the model to learn VI matching more effectively, resulting in enhanced performance in cross-modality Re-ID tasks.
Text-to-Image Person Re-ID. As shown in Tab. 6, ChatReID achieves remarkable gains in text-to-image person Re-ID, especially on the ICFG-PEDES and RSTPReid datasets, where prior performance was relatively low. ChatReID reaches an mAP of 70.5% on ICFG-PEDES, marking a 25.6% improvement, and an mAP of 73.1% on RSTPReid, a 20.2% increase. These significant gains can be attributed to two key factors. First, existing LVLMs already possess a baseline level of text-to-image matching capability, as evidenced in our ablation studies. Second, throughout all three training stages, we utilize a dataset structure based on image-text pairs, which substantially enhances the model’s ability to learn pedestrian-specific text-to-image matching.
Attribute-Based Person Retrieval. To our knowledge, there is currently no straightforward and effective method for the Attribute-Based Person Retrieval task. Existing text-to-image person Re-ID methods primarily focus on retrieving images based on complete textual descriptions rather than specific attributes. To address this gap, we introduced attribute-based retrieval as a training objective in Stage 3 of our model. To assess the effectiveness of our approach, we reorganized the dataset by randomly sampling $N$ images from an attribute-annotated dataset. Our goal was to retrieve images that matched specified attributes, such as clothing or accessories. Experimental results indicate that ChatReID can efficiently retrieve images based on attribute criteria, underscoring its practical value for real-world applications. Additional experimental results are presented in the supplementary material.
Adaptive Multi-Modal Person Retrieval. Considering practical application needs, we design a training task for Adaptive Multi-Modal Person Retrieval. To evaluate the model’s performance on this task, we conducted experiments that included additional textual information as auxiliary input for image-image retrieval tasks. The experimental results demonstrate that ChatReID can efficiently handle person retrieval with multi-modal inputs and that incorporating multi-modal information enhances performance in person Re-ID task. Detailed results and analyses are provided in the supplementary material.
@@ 4.3 Ablation Study
To validate the effectiveness of the three-stage hierarchical progressive tuning strategy, we compare experimental results from different stage combinations. The results are shown in Tab. 5. We first report results from the application of Stage 3 training directly to the pre-trained Qwen2-VL model. The results show that while the pre-trained Qwen2-VL achieves some baseline performance, the results are suboptimal, indicating that current LVLMs possess limited inherent image-image and image-text pedestrian matching capabilities but require specialized training for optimal performance. Performance on the VI-ReID task was particularly low, likely due to the limited proportion of low-illumination data in the training set, which is insufficient for training a robust model. In our experiments, we found that combining only Stage 1 and Stage 3 made it difficult for the model to converge; thus, we did not report those results. However, adding Stage 2 significantly improved model performance, achieving an average mAP of 73.5%, which demonstrates practical utility. This validates the effectiveness of our Stage 2 training design. When Stage 1 training is also included, the model’s performance further improves, reaching SOTA results across ten benchmarks. This final enhancement reinforces the overall effectiveness of our hierarchical three-stage tuning strategy.
5 Conclusion
This paper introduces a novel framework for person re-identification (Re-ID), ChatReID, that leverages the advanced capabilities of large vision-language models (LVLMs). Unlike traditional methods, which are limited by rigid templates and lack of adaptability, our system enables freeform, user-driven queries, offering greater flexibility and natural interaction. By implementing a Hierarchical Progressive Tuning (HPT) strategy, we progressively enhance the model’s ability to achieve fine-grained identity-level retrieval. Our approach, combining Incremental Pre-training, Multi-task Joint Training, and Decoder Post-training , delivers SOTA performance across various Re-ID benchmarks. The system’s VQA inference format simplifies complex Re-ID tasks, making it accessible to non-experts, while its flexibility allows dynamic input combinations and adjustments. Our work establishes a strong foundation for future advancements in person Re-ID, showcasing the potential of LVLMs in practical applications.

### v2

@@ 3.2 Data Engine
In Stage 2, we use seven datasets from various Re-ID settings. We constructed a joint training instruction dataset comprising 3M samples. Given that the current LLM model demonstrates superior text-to-image retrieval capabilities compared to image-to-image retrieval in the person Re-ID task, we augment the data volume for the image-to-image tasks. The two image-to-image tasks account for approximately 25% to 30% of the overall data volume, while the remaining tasks each contribute roughly 10%.
@@
Gallery image sampling The quantity of sampled gallery images for the Many-to-Many Retrieval task adheres to a uniform distribution $N\sim\text{Uniform}(a,b)$. The distribution of the response $P(n)$ is obtained based on inverse transform sampling. This approach allows us to balance the probability that each number falls within different intervals, ensuring that the probability of recording any number eventually is uniform across all numbers. The probability $P(n=k)$ is given by:
@@
In Stage 3, we gather ten widely used benchmarks across four Re-ID settings to create a comprehensive instruction dataset with 500K samples and three specific training objectives. During the training phase, we utilized only the training subsets of these benchmarks, while the inference phase involved evaluation using their respective test sets.
3.3 Details of training and testing
In this section, we detail the technical architecture of ChatReID. The ChatReID framework, illustrated in Figure 3, integrates the structure of Qwen2-VL [39] due to its Naive Dynamic Resolution architecture, which effectively addresses the challenges posed by variable image resolutions in person Re-ID tasks.
In Stage 1, both the encoder and decoder are initialized with a pre-trained VLM to leverage existing language-vision knowledge. We conduct full-parameter training on both the vision encoder and language decoder of Qwen2-VL 2B to enhance the encoder’s focus on pedestrian features, while training the decoder ensures that captions accurately reflect pedestrian attributes. We standardize images with resolutions below 256$\times$128, resizing them to 256$\times$128 while leaving higher-resolution images unchanged. This adjustment leverages our experience with pedestrian tasks, and moderate image compression enhances training efficiency.
In Stage 2, we carry over the encoder from Stage 1 but replace the decoder to continue full-parameter training. The previous decoder’s strong captioning focus could hinder outcomes in this stage, so we use a freshly pre-trained decoder instead. During this stage, we adjust the size of input images to accommodate a large number of gallery images. Images larger than 384$\times$192 were rescaled to 384$\times$192, while images below this threshold were left unaltered.
In Stage 3, we inherit the full structure from Stage 2, training with the encoder frozen. With pedestrian recognition skills already developed, this stage focuses on enhancing the model’s instruction-following capabilities. The input image resolution is preserved without any adjustments.
We rethink the assessment of the person Re-ID problem and formulate a question construction in the style of Visual Question Answering (VQA) utilizing a multi-modal language model from a practical application perspective. The instruction tuning data amounts to 500K.
Person Re-ID. The most common type of query involves providing a text or image query alongside a gallery set, with the goal of identifying images in the gallery that match the query’s identity. However, for current VLMs, retrieving an indefinite-length sequence of images from an indefinite gallery set based on specific criteria is highly challenging. To address this, we transform this complex retrieval task into a multi-turn best-choice problem through engineering enhancements. Specifically, at each step, the model selects the most similar image, making a locally optimal choice, which simplifies the problem and reduces computational complexity.
@@ 3.3 Details of training and testing
Adaptive Multi-Modal Person Retrieval. Finally, in practical scenarios, multi-turn retrieval based on limited query information often results in a set of responses with relatively low accuracy. As retrieval continues, users typically gather more query details to improve subsequent searches. For example, beginning with attribute-based retrieval, if the correct match appears among the initial results, the user can use that match as a new query to refine further results. Alternatively, users can start with image-based retrieval and add textual details to address complexities such as clothing changes or long-term recognition tasks. By integrating multi-modal information, this approach significantly improves both the flexibility and accuracy of person Re-ID, advancing its real-world applicability. To support this, we design a training objective for handling mixed queries with multi-modal inputs.
@@ 4 Experiments
4.1 Experimental Setup
Implementation details are provided in the supplementary.
Datasets. We use the test sets from the datasets employed in Stage 3. The evaluation is conducted using standard metrics in person Re-ID, including Cumulative Match Characteristic (CMC) and Mean Average Precision (mAP). Results are reported in terms of mAP and Rank-1 accuracy.
Evaluation strategy. Evaluating VLMs for person Re-ID in a VQA format presents two primary challenges. First, the large gallery sets in person Re-ID datasets often exceed the token limits of current VLMs. Second, retrieving a sequence of images of variable length based on specific criteria from such vast galleries is inherently difficult for these models. To address these challenges, we implement several engineering optimizations. We first adopt a baseline model (e.g., ResNet-50) to filter the gallery by calculating feature similarity between the query image and gallery images. Only images with similarity above a threshold $\tau$ are retained, significantly reducing the gallery size while maintaining accuracy and improving evaluation efficiency. Next, we reformulate the retrieval task as a multi-turn best-choice problem. Finally, we sample non-overlapping images from the filtered gallery, pairing each selected image with the query and ranking them based on response confidence, ultimately producing the final similarity list.
@@ 4.1 Experimental Setup
Table 2: Comparison of ChatReID with SOTA methods across two cloth-changing Re-ID (CC Re-ID) datasets.
@@ 4.2 Experimental Results
Standard Person Re-ID. As shown in Tab. 1, ChatReID achieves remarkable results across all standard person Re-ID datasets. The most notable improvement is observed on the challenging and large-scale MSMT17 dataset, where ChatReID achieves a 15.1% improvement in mAP. This substantial gain demonstrates the effectiveness of our three-stage tuning strategy, which effectively extracts robust, pedestrian-specific features, thereby improving the model’s resilience in demanding scenarios. Such performance improvement further validates the effectiveness of ChatReID.
Cloth-changing Re-ID (CC Re-ID). The results in Tab. 2 indicate that ChatReID demonstrates significant mAP enhancements over IRM on both the LTCC and PRCC datasets, with increases of 24.3% and 23.3% in mAP, respectively. Based on our analysis, this improvement can be attributed to two main reasons. First, all three tuning stages are based on image-text pairs, where textual information aids the visual encoder learn variations in clothing more effectively. Second, we introduced innovative attribute-based training tasks in Stage 2, which enhance the model’s ability to recognize CC scenarios.
Table 3: Comparison of ChatReID with SOTA methods across two visible-infrared person ReID (VI-ReID) datasets.
@@
Table 4: Ablation Study. Performance comparison of different stage combinations in our hierarchical progressive learning strategy across ten benchmarks. The abbreviation ‘MM01‘ refers to the SYSU-MM01 dataset, ‘CUHK‘ refers to the CUHK-PEDES dataset, and ‘ICFG‘ refers to the ICFG-PEDES dataset.
@@
Market1501 MSMT17 CUHK03 PRCC LTCC MM01 RegDB CUHK ICFG RSTPReid
@@
Visible-infrared person ReID (VI-ReID). As shown in Tab. 3, ChatReID also demonstrates significant improvements on VI-ReID datasets. Specifically, it achieves an 11.8% mAP increase on SYSU-MM01 and a 10.7% mAP increase on RegDB. These gains are primarily attributed to our approach in Stage 2, where we treated the VI-ReID task as an image-image matching task and slightly increased the proportion of VI data in the tuning process.
Text-to-Image Person Re-ID. As shown in Tab. 5, ChatReID achieves remarkable gains in text-to-image person Re-ID, especially on the ICFG-PEDES and RSTPReid datasets, where previous performance was relatively low. ChatReID reaches an mAP of 70.5% on ICFG-PEDES, marking a 25.6% improvement, and an mAP of 73.1% on RSTPReid, a 20.2% increase. These gains are attributed to two factors: existing VLMs have a baseline text-to-image matching capability, and our image-text pair dataset structure enhances person-specific matching across all training stages.
Attribute-Based Person Retrieval. To the best of our knowledge, there is currently no straightforward and effective method for the Attribute-Based Person Retrieval task. Existing text-to-image person Re-ID methods primarily focus on retrieving images based on complete textual descriptions rather than specific attributes. To address this gap, we introduced attribute-based retrieval as a training objective in Stage 3. To assess the effectiveness of our approach, we reorganized the dataset by randomly sampling $N$ images from an attribute-annotated dataset. Our goal was to retrieve images that matched specified attributes, such as clothing or accessories. Experimental results indicate that ChatReID can efficiently retrieve images based on attribute criteria, highlighting its practical value for real-world applications. Additional experimental results are presented in the supplementary material.
Adaptive Multi-Modal Person Retrieval. Considering practical application requirements, we design a training task for Adaptive Multi-Modal Person Retrieval. To evaluate the model’s performance on this task, we conducted experiments that included additional textual information as auxiliary input for image-image retrieval tasks. The experimental results demonstrate that ChatReID can efficiently handle person retrieval with multi-modal inputs. Detailed results and analyses are provided in the supplementary material.
@@ 4.3 Ablation Study
Three-Stage HPT Strategy. To validate the effectiveness of our three-stage HPT strategy, we compare experimental results across different stage combinations, as shown in Tab. 4. Initially, we report the results from directly applying Stage 3 training. While Stage 3 achieves baseline performance, the results are suboptimal, indicating that current VLMs have limited inherent capabilities for image-to-image and image-text pedestrian matching and require specialized training for optimal performance. Our experiments further reveal that combining only Stage 1 and Stage 3 made it difficult for the model to converge, and thus, those results were not reported. However, incorporating Stage 2 into the training process significantly improved the model’s performance, achieving an average mAP of 73.5%. The fine-grained image retrieval tuning in Stage 2 led to substantial improvements, providing strong evidence for the effectiveness of our approach. Finally, when Stage 1 training is also included, the model’s performance is further enhanced, reaching state-of-the-art (SOTA) results across ten benchmarks. This final improvement underscores the overall effectiveness of our hierarchical three-stage tuning strategy.
VLMs Model Size and Performance Trade-off. In our validation experiments, we evaluated VLMs of different sizes under each setting. The results are shown in Tab. 6. As the model size increases, accuracy improves. Both the 2B and 7B models achieve comparable performance. This can be attributed to the fact that the Re-ID task is relatively less complex compared to VLM training scenarios, allowing the 2B model to effectively capture the necessary representations. Given that the 7B model incurs substantially higher computational costs without yielding significant performance gains, we chose the 2B model.
Table 6: Performance on VLMs with different model size. The data in the table represents Rank-1.
#Parameter MSMT17 LTCC RegDB CUHK-PEDES
0.5 B 38.1 27.4 4.3 55.1
2 B 90.1 82.7 96.5 83.8
7 B 91.3 82.5 96.3 87.7
5 Broad Impact and Conclusion
Broad Impact. We utilize publicly available, high-quality academic datasets for our research. These datasets were curated by various institutions across diverse scenarios, which helps mitigate potential biases in the images. It is important to note that, due to privacy policies, we do not use the DukeMTMC [31] dataset. We will release the code, datasets, and models upon acceptance to facilitate further research. To ensure responsible use, we will adhere to a strict application protocol to prevent our work from being used for any unethical or illegal purposes.
Conclusion. This paper introduces ChatReID, a novel framework for person re-identification (Re-ID) that leverages the advanced capabilities of vision-language models (VLMs). Our ChatReID is a text-side-dominated framework for person Re-ID. Different from traditional methods that calculates image feature distance as similarity, ChatReID interprets the task requirements from the input text description and performs similarity inference on the given person images accordingly. By implementing a Hierarchical Progressive Tuning (HPT) strategy, we progressively enhance the model’s ability to achieve fine-grained, identity-level retrieval. The system’s VQA-based inference format simplifies complex Re-ID tasks, making it more accessible to non-experts, while its flexibility allows dynamic input combinations and adjustments. Our work lays a strong foundation for future advancements in person Re-ID, showcasing the potential of LVLMs in practical applications.